a stone is dropped from a minar of height h and it reaches after t seconds on earth. from the same minar if two stones are thrown(one upwards and one downwards) with the same elocity u and they reach the earth surface after t1 and t2 seconds respectiely, then
 
  1. t=t1-t2
  2. t=(t1+t2) /2
  3. t=root of  t1t2
  4. t=t12+t22

Grade:11

3 Answers

Nicho priyatham
8 years ago
C) t= root t1t2
PLZ APPROVE IF CORRECT n exp will b ready 
Nicho priyatham
8 years ago
rember 1 thing 
for any such  problem in kinamitcs u need to 1st u first need to wright the equation 
sf=si+ut +(1/2)gt2
remember this is a vector equation
so here u have 3 cases
  • case1(t)
              it is clear that t=root 2h/g
  • now case 2(t1)
            apply the above equation
          take ground as orisin 
          so sf=0 si=h   u=u a=-g
     so   0=h+ut1+(1/2)gt12  ______________________(1)
  • now case 3(t2)
            appling the  same equation
          so sf=0 si=h   u=-u a=-g
     so   0=h-ut2-(1/2)gt22  ______________________(2)
 
  • now to eliminate u
             multiply t2 with eq1 and add with t1 multiplied with eq2
              t2(eq 1) +  t(eq2)
               and take ( t1 +t2) common and solve to get the answer 
PLZ APPROVE IF useful in case of any doubts ask in ans box
Archita
8 years ago
thanks